OpenStudy (anonymous):

Find the integral using substitution: (2+ ln x)^1/2/(x)

OpenStudy (anonymous):

let u = (2+ln x)

ganeshie8 (ganeshie8):

\(\large \frac{(2+ ln x)^{\frac{1}{2}}}{x}\)

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

What is du?

OpenStudy (anonymous):

Yes, ganeshie8, just like that.

OpenStudy (anonymous):

du= 1/x?

ganeshie8 (ganeshie8):

since we can deal wid ENTIRE thing inside radical at a time, simply sub whatever inside radical as u

OpenStudy (anonymous):

yes...so what form does the integral have?

OpenStudy (anonymous):

"dx" if that is what you are looking for.

OpenStudy (anonymous):

If u = (2+lnx), then du = dx/x...agree?

OpenStudy (anonymous):

dx/x = (1/x) dx

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So the entire integrand called (2 + ln x)^(1/2) dx/x transforms into the form u^(1/2) du....agree?

OpenStudy (anonymous):

because u = 2+ln x and everything else is just du, so you have u^1/2 du

OpenStudy (anonymous):

Yes i got to that step and then I got confused and did not know what to do.

OpenStudy (anonymous):

the integral of u^(1/2) du is just u^(3/2)/(2/3) = 2/3 (u^(3/2) + C Now just substitute (2+ln x) for u

OpenStudy (anonymous):

Oh, I think I over complicated it when I was trying to work through it. Thank you Easyaspi314.

OpenStudy (anonymous):

welcome.